Assume that the population proportion is .55. Compute the standard error of the proportion, σ ( p ¯ ), for sample sizes of 100, 200, 500, and 1000 (to 4 decimals).

Answer by stanbon(75887)

Assume that the population proportion is .55. Compute the standard error of the proportion, σ ( p ¯ ), for sample sizes of 100, 200, 500, and 1000 (to 4 decimals).
------
====
s(phat) = sqrt[0.55*0.45/100] = 0.0497
------
s(phat) = sqrt(0.55*0.45/200) = 0.0352
----------------------
etc.
Cheers,
Stan H.